OpenStudy (anonymous):

(sin^2xcosx)/(cos^2xsinx)

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{sin^2x~cosx}{cos^2x~sinx} }\)

OpenStudy (solomonzelman):

do you see that you can cancel things out ?

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{sin(x)~sin(x)~cos(x)}{cos(x)~cos(x)~sin(x)} }\) \(\Huge\color{blue}{ \sf \frac{sin(x)~\cancel{ sin(x)} ~\cancel{ cos(x)}}{cos(x)~\cancel{ cos(x)}~\cancel{ sin(x)} } }\) \(\Huge\color{blue}{ \sf \frac{sin(x)}{cos(x)} }\) This is equal to one trigonometric function. can you tell me what is it ?

OpenStudy (anonymous):

Tanx